# Negative Sampling in Recommendation: A Survey and Future Directions

### v1

@@ Acknowledgements.
- Liu et al. (2019) Hao Liu, Ting Li, Renjun Hu, Yanjie Fu, Jingjing Gu, and Hui Xiong. 2019. Joint representation learning for multi-modal transportation recommendation. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 33. 1036–1043.
- Liu et al. (2023) Junling Liu, Chao Liu, Peilin Zhou, Renjie Lv, Kang Zhou, and Yan Zhang. 2023. Is ChatGPT a Good Recommender? A Preliminary Study. (2023). http://arxiv.org/abs/2304.10149
@@
- Ma et al. (2022) Yunshan Ma, Yingzhi He, An Zhang, Xiang Wang, and Tat-Seng Chua. 2022. CrossCBR: Cross-view Contrastive Learning for Bundle Recommendation. arXiv preprint arXiv:2206.00242 (2022).
@@
- Pan et al. (2008) Rong Pan, Yunhong Zhou, Bin Cao, Nathan N. Liu, Rajan Lukose, Martin Scholz, and Qiang Yang. 2008. One-Class Collaborative Filtering. In Proceedings of the IEEE International Conference on Data Mining (ICDM).
@@
- Quadrana et al. (2017) Massimo Quadrana, Alexandros Karatzoglou, Balázs Hidasi, and Paolo Cremonesi. 2017. Personalizing session-based recommendations with hierarchical recurrent neural networks. In proceedings of the Eleventh ACM Conference on Recommender Systems. 130–137.
@@
- Rendle et al. (2012) Steffen Rendle, Christoph Freudenthaler, Zeno Gantner, and Lars Schmidt-Thieme. 2012. BPR: Bayesian personalized ranking from implicit feedback. arXiv preprint arXiv:1205.2618 (2012).
- Saito et al. (2020) Yuta Saito, Suguru Yaginuma, Yuta Nishino, Hayato Sakata, and Kazuhide Nakata. 2020. Unbiased recommender learning from missing-not-at-random implicit feedback. In Proceedings of the 13th International Conference on Web Search and Data Mining. 501–509.
@@
- Shi et al. (2023) Wentao Shi, Jiawei Chen, Fuli Feng, Jizhi Zhang, Junkang Wu, Chongming Gao, and Xiangnan He. 2023. On the Theories Behind Hard Negative Sampling for Recommendation. arXiv preprint arXiv:2302.03472 (2023).
@@
- Wang et al. (2019b) Hongwei Wang, Fuzheng Zhang, Miao Zhao, Wenjie Li, Xing Xie, and Minyi Guo. 2019b. Multi-task feature learning for knowledge graph enhanced recommendation. In The world wide web conference. 2000–2010.
@@
- Wang et al. (2019a) Xiang Wang, Xiangnan He, Yixin Cao, Meng Liu, and Tat-Seng Chua. 2019a. Kgat: Knowledge graph attention network for recommendation. In Proceedings of the 25th ACM SIGKDD international conference on knowledge discovery & data mining. 950–958.
@@
- Wang et al. (2024) Yidan Wang, Zhaochun Ren, Weiwei Sun, Jiyuan Yang, Zhixiang Liang, Xin Chen, Ruobing Xie, Su Yan, Xu Zhang, Pengjie Ren, et al. 2024. Enhanced generative recommendation via content and collaboration integration. arXiv preprint arXiv:2403.18480 (2024).
@@
- Wu et al. (2019d) Chuhan Wu, Fangzhao Wu, Mingxiao An, Jianqiang Huang, Yongfeng Huang, and Xing Xie. 2019d. NPA: neural news recommendation with personalized attention. In Proceedings of the 25th ACM SIGKDD international conference on knowledge discovery & data mining. 2576–2584.
@@
- Wu et al. (2023) Likang Wu, Zhi Zheng, Zhaopeng Qiu, Hao Wang, Hongchao Gu, Tingjia Shen, Chuan Qin, Chen Zhu, Hengshu Zhu, Qi Liu, Hui Xiong, and Enhong Chen. 2023. A Survey on Large Language Models for Recommendation. (2023). http://arxiv.org/abs/2305.19860
@@
- Wu et al. (2024) Xuansheng Wu, Huachi Zhou, Yucheng Shi, Wenlin Yao, Xiao Huang, and Ninghao Liu. 2024. Could Small Language Models Serve as Recommenders? Towards Data-centric Cold-start Recommendation. In Proceedings of the ACM on Web Conference. 3566–3575.
- Wu et al. (2019e) Yongkai Wu, Lu Zhang, and Xintao Wu. 2019e. Counterfactual fairness: Unidentification, bound and algorithm. In Proceedings of the twenty-eighth international joint conference on Artificial Intelligence.
- Xi et al. (2023) Yunjia Xi, Weiwen Liu, Jianghao Lin, Jieming Zhu, Bo Chen, Ruiming Tang, Weinan Zhang, Rui Zhang, and Yong Yu. 2023. Towards Open-World Recommendation with Knowledge Augmentation from Large Language Models. (2023). http://arxiv.org/abs/2306.10933
@@
- Yang et al. (2017) Carl Yang, Lanxiao Bai, Chao Zhang, Quan Yuan, and Jiawei Han. 2017. Bridging collaborative filtering and semi-supervised learning: a neural approach for poi recommendation. In Proceedings of the 23rd ACM SIGKDD international conference on knowledge discovery and data mining.
@@
- Zhang et al. (2023) Jizhi Zhang, Keqin Bao, Yang Zhang, Wenjie Wang, Fuli Feng, and Xiangnan He. 2023. Is ChatGPT Fair for Recommendation? Evaluating Fairness in Large Language Model Recommendation. (2023). https://doi.org/10.1145/3604915.3608860
@@
- Zhang et al. (2024) Tao Zhang, Luwei Yang, Zhibo Xiao, Wen Jiang, and Wei Ning. 2024. On Practical Diversified Recommendation with Controllable Category Diversity Framework. In Companion Proceedings of the ACM on Web Conference 2024. 255–263.
@@
- Zheng et al. (2020) Yu Zheng, Chen Gao, Xiang Li, Xiangnan He, Yong Li, and Depeng Jin. 2020. Disentangling user interest and popularity bias for recommendation with causal embedding. arXiv preprint arXiv:2006.11011 (2020), 64.

### v2

@@ Acknowledgements.
- Liu et al. (2023b) Junling Liu, Chao Liu, Peilin Zhou, Renjie Lv, Kang Zhou, and Yan Zhang. 2023b. Is ChatGPT a Good Recommender? A Preliminary Study. (2023). http://arxiv.org/abs/2304.10149
@@
- Liu et al. (2023a) Xiaoyang Liu, Chong Liu, Pinzheng Wang, Rongqin Zheng, Lixin Zhang, Leyu Lin, Zhijun Chen, and Liangliang Fu. 2023a. UFNRec: Utilizing False Negative Samples for Sequential Recommendation. In Proceedings of the 2023 SIAM International Conference on Data Mining (SDM). SIAM, 46–54.
@@
- Ma et al. (2025) Haokai Ma, Yunshan Ma, Ruobing Xie, Lei Meng, Jialie Shen, Xingwu Sun, Zhanhui Kang, and Tat-Seng Chua. 2025. Large Language Model Empowered Recommendation Meets All-domain Continual Pre-Training. arXiv preprint arXiv:2504.08949 (2025).
@@
- Ma et al. (2024c) Haokai Ma, Ruobing Xie, Lei Meng, Yimeng Yang, Xingwu Sun, and Zhanhui Kang. 2024c. SeeDRec: Sememe-based Diffusion for Sequential Recommendation. In Proceedings of the Thirty-Third International Joint Conference on Artificial Intelligence, IJCAI-24. International Joint Conferences on Artificial Intelligence Organization, 2270–2278. https://doi.org/10.24963/ijcai.2024/251
- Ma et al. (2024d) Haokai Ma, Yimeng Yang, Lei Meng, Ruobing Xie, and Xiangxu Meng. 2024d. Multimodal Conditioned Diffusion Model for Recommendation. In Companion Proceedings of the ACM Web Conference 2024 (WWW ’24). 1733–1740. https://doi.org/10.1145/3589335.3651956
- Ma et al. (2022) Yunshan Ma, Yingzhi He, An Zhang, Xiang Wang, and Tat-Seng Chua. 2022. CrossCBR: cross-view contrastive learning for bundle recommendation. In Proceedings of the 28th ACM SIGKDD conference on knowledge discovery and data mining. 1233–1241.
@@
- Meng et al. (2020) Lei Meng, Fuli Feng, Xiangnan He, Xiaoyan Gao, and Tat-Seng Chua. 2020. Heterogeneous fusion of semantic and collaborative information for visually-aware food recommendation. In Proceedings of the 28th ACM international conference on multimedia. 3460–3468.
@@
- Rendle et al. (2009) Steffen Rendle, Christoph Freudenthaler, Zeno Gantner, and Lars Schmidt-Thieme. 2009. BPR: Bayesian personalized ranking from implicit feedback. In Proceedings of the Twenty-Fifth Conference on Uncertainty in Artificial Intelligence (Montreal, Quebec, Canada) (UAI ’09). AUAI Press, Arlington, Virginia, USA, 452–461.
- Saito (2020) Yuta Saito. 2020. Asymmetric tri-training for debiasing missing-not-at-random explicit feedback. In Proceedings of the 43rd International ACM SIGIR Conference on Research and Development in Information Retrieval. 309–318.
- Saito et al. (2020a) Yuta Saito, Suguru Yaginuma, Yuta Nishino, Hayato Sakata, and Kazuhide Nakata. 2020a. Unbiased recommender learning from missing-not-at-random implicit feedback. In Proceedings of the 13th International Conference on Web Search and Data Mining. 501–509.
- Saito et al. (2020b) Yuta Saito, Suguru Yaginuma, Yuta Nishino, Hayato Sakata, and Kazuhide Nakata. 2020b. Unbiased recommender learning from missing-not-at-random implicit feedback. In Proceedings of the 13th international conference on web search and data mining. 501–509.
@@
- Shi et al. (2023) Wentao Shi, Jiawei Chen, Fuli Feng, Jizhi Zhang, Junkang Wu, Chongming Gao, and Xiangnan He. 2023. On the theories behind hard negative sampling for recommendation. In Proceedings of the ACM Web Conference 2023. 812–822.
@@
- Tang et al. (2024b) Hao Tang, Guoshuai Zhao, Jing Gao, and Xueming Qian. 2024b. Personalized Representation With Contrastive Loss for Recommendation Systems. IEEE Trans. Multim. 26 (2024), 2419–2429. https://doi.org/10.1109/TMM.2023.3295740
- Tang et al. (2024a) Jiakai Tang, Sunhao Dai, Zexu Sun, Xu Chen, Jun Xu, Wenhui Yu, Lantao Hu, Peng Jiang, and Han Li. 2024a. Towards Robust Recommendation via Decision Boundary-aware Graph Contrastive Learning. In Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining. 2854–2865.
@@
- Wang et al. (2019) Hongwei Wang, Fuzheng Zhang, Miao Zhao, Wenjie Li, Xing Xie, and Minyi Guo. 2019. Multi-task feature learning for knowledge graph enhanced recommendation. In The world wide web conference. 2000–2010.
@@
- Wang et al. (2024a) Yidan Wang, Zhaochun Ren, Weiwei Sun, Jiyuan Yang, Zhixiang Liang, Xin Chen, Ruobing Xie, Su Yan, Xu Zhang, Pengjie Ren, et al. 2024a. Enhanced generative recommendation via content and collaboration integration. arXiv preprint arXiv:2403.18480 (2024).
@@
- Wang et al. (2024b) Zihan Wang, Yonghui Yang, Le Wu, Richang Hong, and Meng Wang. 2024b. Making Non-overlapping Matters: An Unsupervised Alignment enhanced Cross-Domain Cold-Start Recommendation. IEEE Transactions on Knowledge and Data Engineering (2024).
@@
- Wu et al. (2023a) Bin Wu, Xiangnan He, Le Wu, Xue Zhang, and Yangdong Ye. 2023a. Graph-Augmented Co-Attention Model for Socio-Sequential Recommendation. IEEE Trans. Syst. Man Cybern. Syst. 53, 7 (2023), 4039–4051. https://doi.org/10.1109/TSMC.2023.3242308
- Wu et al. (2023b) Bin Wu, Xiangnan He, Qi Zhang, Meng Wang, and Yangdong Ye. 2023b. GCRec: Graph-Augmented Capsule Network for Next-Item Recommendation. IEEE Trans. Neural Networks Learn. Syst. 34, 12 (2023), 10164–10177. https://doi.org/10.1109/TNNLS.2022.3164982
- Wu and Ye (2020) Bin Wu and Yangdong Ye. 2020. BSPR: Basket-sensitive personalized ranking for product recommendation. Inf. Sci. 541 (2020), 185–206. https://doi.org/10.1016/J.INS.2020.06.046
- Wu et al. (2023c) Bin Wu, Lihong Zhong, and Yangdong Ye. 2023c. Graph-Augmented Social Translation Model for Next-Item Recommendation. IEEE Trans. Ind. Informatics 19, 11 (2023), 10913–10922. https://doi.org/10.1109/TII.2023.3242809
@@
- Wu et al. (2024a) Junkang Wu, Jiawei Chen, Jiancan Wu, Wentao Shi, Jizhi Zhang, and Xiang Wang. 2024a. BSL: Understanding and Improving Softmax Loss for Recommendation. In 40th IEEE International Conference on Data Engineering, ICDE 2024, Utrecht, The Netherlands, May 13-16, 2024. IEEE, 816–830. https://doi.org/10.1109/ICDE60146.2024.00068
@@
- Wu et al. (2024b) Jiancan Wu, Xiang Wang, Xingyu Gao, Jiawei Chen, Hongcheng Fu, and Tianyu Qiu. 2024b. On the Effectiveness of Sampled Softmax Loss for Item Recommendation. ACM Trans. Inf. Syst. 42, 4 (2024), 98:1–98:26. https://doi.org/10.1145/3637061
@@
- Wu et al. (2024c) Likang Wu, Zhi Zheng, Zhaopeng Qiu, Hao Wang, Hongchao Gu, Tingjia Shen, Chuan Qin, Chen Zhu, Hengshu Zhu, Qi Liu, et al. 2024c. A survey on large language models for recommendation. World Wide Web 27, 5 (2024), 60.
@@
- Wu et al. (2024d) Xuansheng Wu, Huachi Zhou, Yucheng Shi, Wenlin Yao, Xiao Huang, and Ninghao Liu. 2024d. Could Small Language Models Serve as Recommenders? Towards Data-centric Cold-start Recommendation. In Proceedings of the ACM on Web Conference. 3566–3575.
- Wu et al. (2019d) Yongkai Wu, Lu Zhang, and Xintao Wu. 2019d. Counterfactual fairness: Unidentification, bound and algorithm. In Proceedings of the twenty-eighth international joint conference on Artificial Intelligence.
- Xi et al. (2024) Yunjia Xi, Weiwen Liu, Jianghao Lin, Xiaoling Cai, Hong Zhu, Jieming Zhu, Bo Chen, Ruiming Tang, Weinan Zhang, and Yong Yu. 2024. Towards open-world recommendation with knowledge augmentation from large language models. In Proceedings of the 18th ACM Conference on Recommender Systems. 12–22.
@@
- Yang et al. (2024) Yonghui Yang, Le Wu, Zihan Wang, Zhuangzhuang He, Richang Hong, and Meng Wang. 2024. Graph bottlenecked social recommendation. In Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining. 3853–3862.
- Yang et al. (2023d) Yonghui Yang, Le Wu, Kun Zhang, Richang Hong, Hailin Zhou, Zhiqiang Zhang, Jun Zhou, and Meng Wang. 2023d. Hyperbolic graph learning for social recommendation. IEEE Transactions on Knowledge and Data Engineering 36, 12 (2023), 8488–8501.
- Yang et al. (2023c) Yonghui Yang, Zhengwei Wu, Le Wu, Kun Zhang, Richang Hong, Zhiqiang Zhang, Jun Zhou, and Meng Wang. 2023c. Generative-contrastive graph learning for recommendation. In Proceedings of the 46th international ACM SIGIR Conference on Research and Development in Information Retrieval. 1117–1126.
@@
- Zhang et al. (2023b) An Zhang, Leheng Sheng, Zhibo Cai, Xiang Wang, and Tat-Seng Chua. 2023b. Empowering Collaborative Filtering with Principled Adversarial Contrastive Loss. In Advances in Neural Information Processing Systems 36: Annual Conference on Neural Information Processing Systems 2023, NeurIPS 2023, New Orleans, LA, USA, December 10 - 16, 2023. http://papers.nips.cc/paper_files/paper/2023/hash/13f1750b825659394a6499399e7637fc-Abstract-Conference.html
- Zhang et al. (2023a) Jizhi Zhang, Keqin Bao, Yang Zhang, Wenjie Wang, Fuli Feng, and Xiangnan He. 2023a. Is ChatGPT Fair for Recommendation? Evaluating Fairness in Large Language Model Recommendation. (2023). https://doi.org/10.1145/3604915.3608860
@@
- Zhang et al. (2024a) Kepu Zhang, Teng Shi, Sunhao Dai, Xiao Zhang, Yinfeng Li, Jing Lu, Xiaoxue Zang, Yang Song, and Jun Xu. 2024a. SAQRec: Aligning Recommender Systems to User Satisfaction via Questionnaire Feedback. In Proceedings of the 33rd ACM International Conference on Information and Knowledge Management. 3165–3175.
@@
- Zhang et al. (2024b) Tao Zhang, Luwei Yang, Zhibo Xiao, Wen Jiang, and Wei Ning. 2024b. On Practical Diversified Recommendation with Controllable Category Diversity Framework. In Companion Proceedings of the ACM on Web Conference 2024. 255–263.
@@
- Zheng et al. (2021) Yu Zheng, Chen Gao, Xiang Li, Xiangnan He, Yong Li, and Depeng Jin. 2021. Disentangling user interest and conformity for recommendation with causal embedding. In Proceedings of the Web Conference 2021. 2980–2991.